OpenStudy (anonymous):

A blimp is hovering 120 metres directly above the finish line of a 200 metre sprint. The angle of depression to the first runner on the home straight is 50° initially, and 8 seconds later, it is 85°. Calculate the distance travelled (to the nearest metre) by the runner in the 8 seconds and his average speed in km/h for that part of the race.

OpenStudy (anonymous):

http://www.twiddla.com/895293 come here

OpenStudy (anonymous):

first find the value of x

OpenStudy (anonymous):

did u got it???

OpenStudy (campbell_st):

draw a picture  find the length of b and then a the distance travelled in 8 seconds is b - a tan 50 = 120/b tan85 = 120/a b = 120/tan 50 a = 120/tan85 this should be a good start to the question